# Locality Alignment Improves Vision-Language Models

### v1

@@ 4.1 Probing benchmark
A natural task to test if a ViT encodes local image semantics is semantic segmentation (Long et al., 2015). However, this is a pixel-level classification problem, and the most performant approaches for ViTs require fully fine-tuning the backbone (Li et al., 2022c, Chen et al., 2022b, Fang et al., 2023b), sometimes with windowed self-attention (Li et al., 2022b). We want to test how well a ViT captures local semantics without task-specific fine-tuning, so we simplify the problem by casting it as a patch-level multi-label classification problem and keep the backbone frozen. Specifically, we create a small output head on top of the ViT’s output representation, and we train it to predict the union of labels in each patch using a binary cross-entropy (BCE) loss. We implement this approach with MSCOCO (Lin et al., 2014), but we can also use other datasets like Ade20k (Zhou et al., 2019).
@@ 4.2 Ablating MaskEmbed design choices
Our first usage of the probing benchmark is to explore several design choices for MaskEmbed. There are certain hyperparameters that we did not fully specify in Section 3.2, including the choice of reconstruction target and mask distribution, and we also want to test the importance of data augmentations, training duration and data diversity (IN1k vs. IN21k). We consider two pre-trained models for these experiments, IN1k ViT-B/16 and CLIP ViT-B/16 (Dosovitskiy et al., 2020, Radford et al., 2021), and we conduct a series of ablations to investigate these implementation choices.
We report the full results of our ablations in Appendix B, but we describe our main findings here that inform settings for our later runs. Reconstruction target: we observe that reconstructing the [CLS] token improves local probing performance, but not as much as reconstructing the entire embedding sequence from the second-to-last layer; this is expected, and we adopt this choice for the rest of our experiments. Mask sampling: we find that multiple mask distributions are effective, including the block masking approach from BEiT (Bao et al., 2021). We adopt an unstructured mask whose cardinality is sampled uniformly at random, and we additionally train with the complement of the mask and a mask that preserves all patches at each iteration. [^2] In our notation this corresponds to $p(m)=1/\binom{n}{|m|}(n+1)$, and at each step we calculate the reconstruction loss for three masks: $m\sim p(m)$, $1-m$ and ${\bm{1}}$. Data augmentations: we observe that strong augmentations like Mixup, CutMix and AutoAugment are not necessary for improved performance (Zhang et al., 2017, Yun et al., 2019, Cubuk et al., 2018), and we use a simple random crop for our main runs. Decoder size: performance is not overly sensitive to the decoder size, so we adopt a simple two-layer transformer. Training data: we find that local probing performance improves within just 2 IN1k epochs, and that we can get strong improvements in under 50 epochs. We also find that training with the more diverse IN21k is important for CLIP ViT-B/16, which is pre-trained with more diverse data and can degrade when fine-tuned for too long with IN1k. For our remaining runs we train all models with IN21k for 5 epochs, which is equivalent to roughly 60k gradient steps with batch size 1024. Notably, this is less than 1% of pre-training cost for CLIP and SigLIP (Radford et al., 2021, Zhai et al., 2023b), so the marginal cost of locality alignment is low.
@@ 4.3 Comparison with pre-trained models
We now perform experiments to verify that MaskEmbed improves local feature extraction for a range of pre-trained models. We consider ViTs trained with multiple forms of image-level supervision, including IN1k classifiers (Dosovitskiy et al., 2020), CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023b), other language-supervised models (OpenCLIP, DFN, EVA02; Cherti et al. 2023, Fang et al. 2023a, 2024) and MoCo v3 (Chen et al., 2021). Not all of these models are relevant for high-performance VLMs (Tong et al., 2024a), but locality alignment should work for any model pre-trained with image-level supervision. We use the settings determined in our ablation study, which include reconstructing the teacher’s entire embedding sequence and training with IN21k for 5 epochs. Other details on our MaskEmbed hyperparameters are described in Appendix C.
Overall, we find that MaskEmbed reliably improves local probing performance for all these models, and in many cases even improves their global probing performance. Figure 4 (left) shows the local probing accuracy for IN1k models across different scales, where we find that performance improves for all models except the low-capacity ViT-T: locality alignment boosts the ViT-B’s performance to roughly that of the next model scale, and provides a similar absolute improvement for the ViT-L. Next, Figure 4 (right) shows results for a range of models, including three CLIP and three SigLIP backbones, all of which improve substantially. Notably, the two strongest backbones for VLMs show clear improvements (CLIP ViT-L @ 336px and SigLIP SO400M @ 384px), suggesting that the challenge of learning local semantics is not solved merely with scale. Figure 3 shows qualitative examples from CLIP ViT-L @ 336px, demonstrating how MaskEmbed helps identify where each object occurs in the image. Appendix B shows results for the remaining models, all of which show similarly large improvements (OpenCLIP, DFN, EVA02, MoCo v3); we even find that locality alignment can improve probing performance for some densely supervised models, including BEiT and BEiTv2 (Bao et al., 2021, Peng et al., 2022).
@@
Finally, we perform a comparison with CLIPSelf (Wu et al., 2024). This method uses a similar objective and reconstructs cropped views using cropped ViT features, but it reconstructs CLIP’s [CLS] token by simply average-pooling embeddings within each crop window. We test this method in Table 1, where we find that it in fact degrades CLIP’s probing performance. We suspect that the main issue is not crops but the use of a weak decoder (i.e., averaging features within the crop), and we verify that MaskEmbed also degrades performance when we use this approach to reconstruct the [CLS] token. Our main version of MaskEmbed proves to be much more effective, although unlike CLIPSelf it does not preserve CLIP’s zero-shot classification abilities.
@@ 5 Vision-Language Experiments
Experimental setup. We train VLMs using the Prismatic library and training recipe (Karamcheti et al., 2024). Images are turned into embedding sequences by the ViT (Liu et al., 2023c), projected into the LM embedding space by an adapter module, concatenated with text token embeddings, and processed by the LM. We train in a single stage with the ViT frozen, following Karamcheti et al. (2024). Our experiments focus on two high-resolution vision backbones, CLIP ViT-L @ 336px and SigLIP SO400M @ 384px (Radford et al., 2021, Zhai et al., 2023b, Alabdulmohsin et al., 2023), which respectively have 306M and 400M parameters and represent images with 577 and 729 tokens. For our LM backbone we use Llama-2 7B Base (Touvron et al., 2023), which was found to outperform the instruction-tuned Vicuña 7B (Zheng et al., 2023) by Karamcheti et al. (2024).
For our training dataset, we use the Llava-1.5 data mixture (Liu et al., 2024) that contains 665k examples, and which consists of synthetic instruction completions (Liu et al., 2023c), existing vision-language datasets (e.g., GQA, TextCaps; Hudson and Manning 2019, Sidorov et al. 2020) and a collection of language-only data (ShareGPT, 2023). We also experiment with an extended data mixture considered by Karamcheti et al. (2024), which adds LVIS-Instruct-4V (Wang et al., 2023a) and LRV-Instruct (Liu et al., 2023b) for an additional 570k examples. We provide more details on the training data in Appendix D, and all models are trained for two epochs.
Evaluations. We use a suite of standardized benchmarks considered by Karamcheti et al. (2024). Those that involve spatial understanding and fine-grained features include object localization (RefCOCO, OCID-Ref; Kazemzadeh et al. 2014, Wang et al. 2021), counting (TallyQA; Acharya et al. 2019), relational question-answering (VSR; Liu et al. 2023a), chart understanding (AI2D; Kembhavi et al. 2016) and text comprehension (TextVQA; Singh et al. 2019). We also show results for holistic question-answering (VQAv2, VizWiz; Goyal et al. 2017, Bigham et al. 2010) and object hallucination (POPE; Li et al. 2023c), which are not as closely related to spatial understanding. We provide more details on our suite of benchmarks in Appendix D.
@@ 5.1 Results
For VLMs trained with standard backbones, we follow the exact training recipe from Karamcheti et al. (2024). But for those trained with locality-aligned backbones, we find that one small architecture change is necessary to achieve these performance improvements: rather than using the standard MLP vision-language adapter (Liu et al., 2024), we use the trained decoder module from MaskEmbed as an adapter (see Section 3.2). This unlocks robust performance improvements consistent with our probing results in Section 4.3, whereas using a MLP adapter applied to the fine-tuned embeddings slightly hurts performance (see ablations in Appendix D). We reason that this is because information is compressed into a space that is difficult to use compared to the text-aligned CLIP and SigLIP spaces, and that the decoder helps resolve this for the LM. Overall, the modified adapter adds negligible compute overhead and is a simple change to yield improved spatial understanding.
In Appendix D, we also show a comparison with an alternative approach to improving spatial understanding: fusing features from a second backbone, specifically DINOv2 (Oquab et al., 2023), following the implementation from Karamcheti et al. (2024). We find that both methods improve spatial understanding benchmarks like RefCOCO and TallyQA, with feature fusion in some cases leading to larger gains. However, we also observe that feature fusion can degrade the model in other ways that do not occur with locality alignment, including holistic question-answering (VizWiz) and text comprehension (TextVQA) – likely because text is not prominent in DINOv2’s pre-training. We leave to future work a careful study of how to compose locality alignment with feature fusion, as well as other ideas like combining multi-crop features (Liu et al., 2024, Xu et al., 2024b), increasing image resolution (Bai et al., 2023) and utilizing prefix attention in the LM (Beyer et al., 2024).
@@ 6 Discussion
Our main contributions in this work are proposing locality alignment as a post-training stage for ViTs, investigating a specific implementation with MaskEmbed, and demonstrating improvements in local feature extraction and VLM performance (Sections 4 and 5). We find that fixing a vision backbone’s local feature extraction can be done relatively efficiently using only self-supervision, and that this is effective for many models trained with image-level objectives. Most notably, locality alignment boosts performance for VLMs trained with high-resolution CLIP and SigLIP backbones.
One limitation of our work is that we focus on a single VLM training approach – the Llava-style patches-as-tokens architecture (Liu et al., 2023c) and the specific Prismatic recipe of training in a single stage with the ViT frozen (Karamcheti et al., 2024). The benefits of locality alignment may change with end-to-end fine-tuning, but we did not explore this because it is unhelpful with our quantity of multi-modal training data (Karamcheti et al., 2024). An important direction for future work is to test locality alignment in other VLM training approaches, with larger LMs, and to evaluate how it composes with other techniques that enhance visual features.
As other directions for future work, we speculate that locality alignment may yield larger gains when training for longer with more diverse data (e.g., DataComp; Gadre et al. 2023). It may also be possible to iteratively learn from stronger teacher models learned during locality alignment, similar to the momentum encoding approach in data2vec (Baevski et al., 2022). Next, because we observe significant gains for large and high-resolution backbones, an exciting next step is to locality-align native-resolution ViTs (Dehghani et al., 2023b): these offer the potential to capture fine-grained details in large images, but due to their large token counts are at higher risk of mixing information across locations and losing local semantics. And finally, because MaskEmbed can be understood as leveraging synthetic data for large-scale dense supervision, it may be possible to adapt our approach for end-to-end vision-language training and incorporate it into the pre-training data mixture for next-generation models like Chameleon (Chameleon Team, 2024).
@@ Appendix A Extended Related Work
This section provides an extended discussion of related work, including our proposal’s connection to knowledge distillation and its differences with existing pre-training and distillation approaches.
@@
Other ViT pre-training methods. The main text mentions a number of strongly supervised, language-supervised and self-supervised pre-training methods (see Section 2). We add to list this several more self-supervised methods including iBOT (Zhou et al., 2021), DINOv2 (Oquab et al., 2023), MoCo (Chen et al., 2021), CISSL/DISSL (Dubois et al., 2022), and pretext tasks like jigsaw puzzle solving (Noroozi and Favaro, 2016) and rotation prediction (Gidaris et al., 2018). Beyond these works that develop new objectives, other works explore combinations of multiple objectives (Mu et al., 2022, Kim et al., 2023, Dong et al., 2023, Chen et al., 2024), e.g., CLIP combined with SimCLR (Chen et al., 2020b) or MAE (He et al., 2022). Other works combine pre-training with distillation from strong teacher models (Sameni et al., 2024). Our proposal for a locality alignment stage composes with any pre-training approach, but it is most applicable to those that provide large-scale, semantically rich supervision without any localization (e.g., CLIP). Our locality alignment post-training stage removes the need to augment such objectives with either a secondary objective to learn localized semantics.
@@
Knowledge distillation. Knowledge distillation is a technique to train small models that imitate larger ones (Hinton et al., 2015) that works across many machine learning problems (Sanh et al., 2019, Taori et al., 2023). Deviating from this motivation, some works have adopted versions of distillation for self-supervised learning (Caron et al., 2021, Baevski et al., 2022), and others use it for masked image modeling (Peng et al., 2022, Fang et al., 2023b) or to learn models that handle missing information for better interpretability (Frye et al., 2020, Jethani et al., 2021, Jain et al., 2022). MaskEmbed is a form of self-distillation because we reconstruct augmented teacher views, similar to works like Consistent Teaching (Beyer et al., 2022) and ReLabel (Yun et al., 2021). However, our use of masking at the embedding layer is a key difference from these approaches that enables MaskEmbed to learn localized patch semantics.
@@
Comparison with existing approaches. In Table 2, we compare MaskEmbed to existing methods that use various combinations of masked prediction, dense supervision and knowledge distillation. MaskEmbed is unique in its use of dual masking for both the student and teacher, because most methods only perform masking for the student model. Unlike other densely supervised methods, especially masked image modeling methods like MAE, BEiT and MaskFeat (He et al., 2022, Bao et al., 2021, Wei et al., 2022), we do not adopt single labels for each patch: MaskEmbed is the only method in Table 2 that supervises student predictions by decoding arbitrarily masked patch embeddings to reconstruct mask-dependent labels. Overall, MaskEmbed has important differences from prior works that enable learning rich localized semantics from a pre-trained teacher model.
@@
Table 2: Comparison to methods involving combinations of masked prediction, dense supervision and knowledge distillation. Unlike some previous works, we do not adopt single labels for each patch. Unlike previous works, we perform student masking on patch embeddings rather than raw pixels.
Labels Dense Supervision Teacher Masking Student Masking
MAE (He et al., 2022) Raw pixels
MaskFeat (Wei et al., 2022) HOG features
BEiT (Bao et al., 2021) dVAE
BEiTv2 (Peng et al., 2022) Pre-trained model
EVA (Fang et al., 2023b) Pre-trained model
data2vec (Baevski et al., 2022) Momentum encoder
FLIP (Li et al., 2023b) Image captions
CLIPA (Li et al., 2023a) Image captions
Masked Surrogate (Frye et al., 2020) Pre-trained model
Token Labeling (Jiang et al., 2021) Pre-trained model
MaskEmbed (Ours) Pre-trained model

### v2

@@ 4.1 Probing benchmark
A natural task to test whether a ViT encodes local image semantics is semantic segmentation (Long et al., 2015). However, this is a pixel-level classification problem, and the most performant approaches for ViTs require fully fine-tuning the backbone (Li et al., 2022c; Chen et al., 2022b; Fang et al., 2023b), sometimes with windowed self-attention (Li et al., 2022b). We want to test how well a ViT captures local semantics without task-specific fine-tuning, so we simplify the problem by casting it as a patch-level multi-label classification problem and keep the backbone frozen. Specifically, we create a small output head on top of the ViT’s output embeddings, and we train it to predict the union of labels in each patch using a binary cross-entropy (BCE) loss. We implement this approach with MSCOCO (Lin et al., 2014), but we can also use other datasets like Ade20k (Zhou et al., 2019).
@@ 4.2 Ablating MaskEmbed design choices
Our first usage of the probing benchmark is to explore several design choices for MaskEmbed. There are certain hyperparameters that we did not fully specify in Section 3.2, including the choice of reconstruction target and mask distribution, and we also want to test the importance of data augmentations, training duration and data diversity (IN1k vs. IN21k). We consider two pre-trained models for these experiments, IN1k ViT-B/16 and CLIP ViT-B/16 (Dosovitskiy et al., 2020; Radford et al., 2021), and we conduct a series of ablations to investigate these implementation choices.
We report the full results of our ablations in Appendix B, but we describe our main findings here that inform settings for our later runs. Reconstruction target: we observe that reconstructing the [CLS] token improves local probing performance, but not as much as reconstructing the entire embedding sequence from the second-to-last layer; this is expected, and we adopt this choice for the rest of our experiments. Mask sampling: we find that multiple mask distributions are effective, including the block masking approach from BEiT (Bao et al., 2021). We adopt an unstructured mask whose cardinality is sampled uniformly at random, and we additionally train with the complement of the mask and a mask that preserves all patches at each iteration. [^3] In our notation this corresponds to $p(m)=1/\binom{n}{|m|}(n+1)$, and at each step we calculate the reconstruction loss for three masks: $m\sim p(m)$, $1-m$ and ${\bm{1}}$. Data augmentations: we observe that strong augmentations like Mixup, CutMix and AutoAugment are not necessary (Zhang et al., 2017; Yun et al., 2019; Cubuk et al., 2018), and we use a simple random crop for our main runs. Decoder size: performance is not very sensitive to the decoder size, so we adopt a simple two-layer transformer. Training data: we find that local probing performance improves within just 2 IN1k epochs, and that we can get strong improvements in under 50 epochs. We also find that training with the more diverse IN21k is important for CLIP ViT-B/16, which is pre-trained with more diverse data and can degrade when fine-tuned for too long with IN1k. For our remaining runs we train all models with IN21k for 5 epochs, which is equivalent to roughly 60k gradient steps with batch size 1024. Notably, this is less than 1% of pre-training cost for CLIP and SigLIP (Radford et al., 2021; Zhai et al., 2023b), so the marginal cost of locality alignment is low.
@@ 4.3 Comparison with pre-trained models
We now perform experiments to verify that MaskEmbed improves local feature extraction for a range of pre-trained models. We consider ViTs trained with multiple forms of image-level supervision, including IN1k classifiers (Dosovitskiy et al., 2020), CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023b), other language-supervised models (OpenCLIP, DFN, EVA02; Cherti et al. 2023; Fang et al. 2023a; 2024) and MoCo v3 (Chen et al., 2021). Not all of these models are relevant for high-performance VLMs (Tong et al., 2024a), but we aim to test whether locality alignment works for any model pre-trained with image-level supervision. We use the settings determined in our ablation study, which include reconstructing the teacher’s entire embedding sequence and training with IN21k for 5 epochs. Other details on our MaskEmbed hyperparameters are described in Appendix D.
Overall, we find that MaskEmbed reliably improves local probing performance for all these models, and in many cases even improves their global probing performance. Figure 4 (left) shows the local probing accuracy for IN1k models across different scales, where we find that performance improves for all models except the low-capacity ViT-T: locality alignment boosts the ViT-B’s performance to roughly that of the next model scale, and provides a similar absolute improvement for the ViT-L. Next, Figure 4 (right) shows results for a range of models, including three CLIP and three SigLIP backbones, all of which improve substantially. Notably, the two strongest backbones for VLMs show clear improvements (CLIP ViT-L @ 336px and SigLIP SO400M @ 384px), suggesting that the challenge of learning local semantics is not solved merely with scale, but is significantly improved by locality alignment. Figure 3 shows qualitative examples from CLIP ViT-L @ 336px, demonstrating how MaskEmbed helps identify where each object occurs in the image. Appendix B shows results for the remaining models, all of which show similarly large improvements (OpenCLIP, DFN, EVA02, MoCo v3); we find that locality alignment can even improve probing performance for some densely supervised models, including BEiT and BEiTv2 (Bao et al., 2021; Peng et al., 2022). In addition, we corroborate these results by showing that locality alignment improves IN1k classification accuracy, which represents a more challenging global image understanding task (see Appendix C).
@@
Finally, we perform a comparison with CLIPSelf (Wu et al., 2024). This method uses a similar objective and reconstructs cropped views using cropped ViT features, but it reconstructs CLIP’s [CLS] token by simply average-pooling embeddings within each crop window. We test this method in Table 1, where we find that it in fact degrades CLIP’s probing performance. We suspect that the main issue is not crops but the use of a weak decoder (i.e., averaging features within the crop), and we verify that MaskEmbed also degrades performance when we use this approach to reconstruct the [CLS] token (averaging unmasked embeddings rather than passing them to a learned decoder). Our main version of MaskEmbed proves to be much more effective, although unlike CLIPSelf it does not preserve CLIP’s zero-shot classification abilities.
@@ 5 Vision-Language Experiments
Experimental setup. We train VLMs using the Prismatic library and training recipe (Karamcheti et al., 2024). Images are turned into embedding sequences by the ViT (Liu et al., 2023c), projected into the LM embedding space by an adapter module, concatenated with text token embeddings, and processed by the LM. We train in a single stage with the ViT frozen, following Karamcheti et al. (2024). Our experiments focus on two high-resolution vision backbones, CLIP ViT-L @ 336px and SigLIP SO400M @ 384px (Radford et al., 2021; Zhai et al., 2023b; Alabdulmohsin et al., 2023), which respectively have 306M and 400M parameters and represent images with 577 and 729 tokens. For our LM backbone we use Llama-2 7B Base (Touvron et al., 2023), which was found to outperform the instruction-tuned Vicuña 7B (Zheng et al., 2023) by Karamcheti et al. (2024).
@@
For our training dataset, we use the Llava-1.5 data mixture (Liu et al., 2024) that contains 665k examples, and which consists of synthetic instruction completions (Liu et al., 2023c), existing vision-language datasets (e.g., GQA, TextCaps; Hudson & Manning 2019; Sidorov et al. 2020) and a collection of language-only data (ShareGPT, 2023). We also experiment with an extended data mixture considered by Karamcheti et al. (2024), which adds LVIS-Instruct-4V (Wang et al., 2023a) and LRV-Instruct (Liu et al., 2023b) for an additional 570k examples. We provide more details on the training data in Appendix E, and all models are trained for two epochs.
Evaluations. We use a suite of standardized benchmarks considered by Karamcheti et al. (2024). Benchmarks that involve spatial understanding and fine-grained features include object localization (RefCOCO, OCID-Ref; Kazemzadeh et al. 2014; Wang et al. 2021), counting (TallyQA; Acharya et al. 2019), relational question-answering (VSR; Liu et al. 2023a), chart understanding (AI2D; Kembhavi et al. 2016) and text comprehension (TextVQA; Singh et al. 2019). We also show results for holistic question-answering (VQAv2, VizWiz; Goyal et al. 2017; Bigham et al. 2010) and object hallucination (POPE; Li et al. 2023c), which are not as closely related to spatial understanding. We provide more details on our suite of benchmarks in Appendix E.
@@ 5.1 Results
For VLMs trained with standard backbones, we follow the exact training recipe from Karamcheti et al. (2024). But for those trained with locality-aligned backbones, we find that one small architecture change is necessary to achieve these performance improvements: rather than using the standard MLP vision-language adapter (Liu et al., 2024), we use the trained decoder module from MaskEmbed as an adapter (see Section 3.2). This unlocks robust performance improvements consistent with our probing results in Section 4.3, whereas using a MLP adapter applied to the fine-tuned embeddings slightly hurts performance (see ablations in Appendix E). We reason that this is because information is compressed into a space that is difficult to use compared to the text-aligned CLIP and SigLIP spaces, and that the decoder helps resolve this for the LM. Overall, the modified adapter adds negligible compute overhead and is a simple change to yield improved spatial understanding.
In Appendix E, we also show a comparison with an alternative approach to improving spatial understanding: fusing features from a second backbone, specifically DINOv2 (Oquab et al., 2023), following the implementation from Karamcheti et al. (2024). We find that both methods improve spatial understanding benchmarks like RefCOCO and TallyQA, with feature fusion in some cases leading to larger gains. However, we also observe that feature fusion can degrade the model in other ways that do not occur with locality alignment, including holistic question-answering (VizWiz) and text comprehension (TextVQA) – likely because text is not prominent in DINOv2’s pre-training. We leave to future work a careful study of how to compose locality alignment with feature fusion, as well as other ideas like combining multi-crop features (Liu et al., 2024; Xu et al., 2024b), increasing image resolution (Bai et al., 2023) and utilizing prefix attention in the LM (Beyer et al., 2024).
@@ 6 Discussion
Our main contributions in this work are proposing locality alignment as a post-training stage for ViTs, investigating a specific implementation with MaskEmbed, and demonstrating improvements in local feature extraction and VLM performance. We find that local feature extraction can be improved using only self-supervision, and that this is effective for many models trained with image-level objectives. Most notably, locality alignment boosts performance for VLMs that adopt the high-resolution CLIP and SigLIP backbones, which are widely used in recent works.
One limitation of our work is that we focus on a single VLM training approach – the Llava-style patches-as-tokens architecture (Liu et al., 2023c), and the specific Prismatic recipe of training in a single stage with the ViT frozen (Karamcheti et al., 2024). The benefits of locality alignment may change with end-to-end fine-tuning, but we did not explore this because it is unhelpful with our amount of multi-modal training data (Karamcheti et al., 2024). An important direction for future work is to test locality alignment in other VLM training approaches, with larger LMs, and to evaluate how it composes with other techniques that enhance visual features.
As other directions for future work, we speculate that locality alignment may yield larger gains when training for longer with more diverse data (e.g., DataComp; Gadre et al. 2023). Next, because we observe significant gains for large and high-resolution backbones, an exciting direction is to locality-align native-resolution ViTs (Dehghani et al., 2023b): these offer the potential to capture fine-grained details in large images, but due to their large token counts are at higher risk of mixing information across locations and losing local semantics. And finally, because MaskEmbed can be understood as leveraging synthetic data for large-scale dense supervision, it may be possible to adapt our approach for end-to-end vision-language training and incorporate it into the pre-training data mixture for VLMs like Chameleon (Chameleon Team, 2024), and even vision encoder-free architectures like Fuyu (Bavishi et al., 2023).
@@ Appendix A Extended Related Work
This section provides an extended discussion of related work, including our proposal’s connections and differences with existing pre-training and distillation approaches.
@@
Other ViT pre-training methods. The main text mentions a number of strongly supervised, language-supervised and self-supervised pre-training methods (see Section 2). We add to list this several more self-supervised methods including iBOT (Zhou et al., 2021), DINOv2 (Oquab et al., 2023), MoCo (Chen et al., 2021), CISSL/DISSL (Dubois et al., 2022), and pretext tasks like jigsaw puzzle solving (Noroozi & Favaro, 2016) and rotation prediction (Gidaris et al., 2018). Beyond these works that develop new objectives, other works explore combinations of multiple objectives (Mu et al., 2022; Kim et al., 2023; Dong et al., 2023; Chen et al., 2024), e.g., CLIP combined with SimCLR (Chen et al., 2020b) or CLIP combined with MAE (He et al., 2022). Other works combine pre-training with distillation from strong teacher models (Sameni et al., 2024). Compared to these works, locality alignment 1) relies on self-supervision instead of distilling from other strong models, 2) removes the need for augmenting pre-training objectives with secondary objectives to learn localized semantics.
@@
Knowledge distillation. Knowledge distillation is a technique to train small models that imitate larger ones (Hinton et al., 2015) that works across many machine learning problems (Sanh et al., 2019; Taori et al., 2023). Deviating from the original motivation, some works have adopted versions of distillation for self-supervised learning (Caron et al., 2021; Baevski et al., 2022), and others use it for masked image modeling (Peng et al., 2022; Fang et al., 2023b) or to learn models that handle missing information for better interpretability (Frye et al., 2020; Jethani et al., 2021; Jain et al., 2022). MaskEmbed is a form of self-distillation because we reconstruct augmented teacher views, similar to works like Consistent Teaching (Beyer et al., 2022) and ReLabel (Yun et al., 2021). However, our use of masking at the embedding layer is a key difference that enables MaskEmbed to learn localized patch semantics.
@@
Comparison with other existing approaches. In Table 2, we compare MaskEmbed to existing methods that use various combinations of masked prediction, dense supervision and knowledge distillation. MaskEmbed differs in its use of dual masking for both the student and teacher, because most methods only perform masking for the student model. Unlike other densely supervised methods, especially masked image modeling methods like MAE, BEiT and MaskFeat (He et al., 2022; Bao et al., 2021; Wei et al., 2022), we do not adopt single labels for each patch: MaskEmbed is the only method in Table 2 that supervises student predictions by decoding arbitrarily masked patch embeddings to reconstruct mask-dependent labels. Overall, MaskEmbed has important differences from prior works that enable learning rich localized semantics from a pre-trained teacher model.
@@
Table 2: Comparison to methods involving combinations of masked prediction, dense supervision and knowledge distillation. Unlike some previous works, we do not adopt single labels for each patch but instead let them change as a function of the mask. Unlike previous works, we perform student masking on patch embeddings rather than raw pixels.
Labels Dense Supervision Teacher Masking Student Masking MAE (He et al., 2022) Raw pixels MaskFeat (Wei et al., 2022) HOG features BEiT (Bao et al., 2021) dVAE BEiTv2 (Peng et al., 2022) Pre-trained model EVA (Fang et al., 2023b) Pre-trained model data2vec (Baevski et al., 2022) Momentum encoder FLIP (Li et al., 2023b) Image captions CLIPA (Li et al., 2023a) Image captions Masked Surrogate (Frye et al., 2020) Pre-trained model Token Labeling (Jiang et al., 2021) Pre-trained model MaskEmbed (Ours) Pre-trained model